# A Mathematical Theory for Learning Semantic Languages by Abstract Learners

### v1

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated the emergence of capabilities (learned skills) when the number of system parameters and the size of training data surpass certain thresholds. The exact mechanisms behind such phenomena are not fully understood and remain a topic of active research. Inspired by the skill-text bipartite graph model presented in [1] for modeling semantic language, we develop a mathematical theory to explain the emergence of learned skills, taking the learning (or training) process into account. Our approach models the learning process for skills in the skill-text bipartite graph as an iterative decoding process in Low-Density Parity Check (LDPC) codes and Irregular Repetition Slotted ALOHA (IRSA). Using density evolution analysis, we demonstrate the emergence of learned skills when the ratio of the size of training texts to the number of skills exceeds a certain threshold. Our analysis also yields a scaling law for testing errors relative to the size of training texts. Upon completion of the training, we propose a method for semantic compression and discuss its application in semantic communication.
@@ I Introduction
In the recent era of natural language processing (NLP), the evolution of large language models (LLMs), such as GPT-4 [2] and Gemini [3], has greatly impacted people’s daily lives. There is a growing consensus that enhancing language model performance and sample efficiency across a broad spectrum of downstream NLP tasks is closely linked to scaling up these models. This scaling involves increasing both the size of the training data and the number of model parameters, as discussed in [4, 5].
The relationship between the scale of an LLM and its performance can often be quantitatively forecasted using scaling laws, a concept explored in-depth in [6, 7]. Despite this, not all downstream tasks exhibit continuous performance improvement with increased scale. This deviation from the expected trend, where certain tasks do not conform to predictive models, is highlighted in [8].
@@
Further exploration in recent studies [9, 10, 2] demonstrates, through various numerical examples, that LLMs manifest emergent capabilities absent in their smaller counterparts. There is a notable enhancement in system performance once a certain critical scale threshold is surpassed. This phenomenon aligns with the concept of a phase transition as known in network science, a topic elaborately discussed in the book [11]. However, as pointed out in [9], the exact mechanisms behind the development of these emergent abilities in large language models are still not fully understood and remain a topic of active research.
@@
In the study presented by [12], a simple explanation is offered for the phase transition phenomenon observed in LLMs based on the concept of list decoders. The approach in [12] models an LLM as a sequence-to-sequence random function over a certain token space with $M$ possible tokens. It was shown that the expected number of erroneous sequences in the list decoder can be bounded by a constant if $M\epsilon<1$, where $\epsilon$ is the false alarm probability. Since transformer-based LLMs with more parameters and extensive training can memorize more patterns [13, 14], they are more likely to reduce the false alarm probability $\epsilon$ below the percolation threshold of $1/M$. However, the list decoder approach does not explain how the capabilities of language skills emerge from scaling up the number of parameters and the size of training data.
In [1], a novel random graph approach was proposed to explain the scaling law discussed in [7]. A language is modeled by a random bipartite graph with skills on one side and texts (originally termed text-pieces in [1]) on the other side. An edge between a text and a skill indicates that understanding the text requires the skill. A text can be comprehended if all the skills connected to it are acquired.
Inspired by the random bipartite graph approach in [1], we take one step further by incorporating learning (or training). We summarize our contributions as follows:
@@
(i) A mathematical theory for the emergence of learned skills: Specifically, we model the learning process for the skills in the skill-text bipartite graph as an iterative decoding process, similar to the analysis in Low-Density Parity Check (LDPC codes) [15, 16, 17] and Irregular Repetition Slotted ALOHA (IRSA) [18, 19, 20, 21, 22, 23]. When a text is presented to a learner, a skill in that text might be learned with a certain probability (specific definitions for learners are given in Section III). By repeatedly presenting a large number of training texts to a learner, a fraction of skills can be learned. Following the density evolution analysis for the iterative decoding process (in the asymptotic regime of a large number of skills) [24, 25, 26], we show the emergence of learned skills when the ratio $R$ of the number of training texts to the number of skills exceeds a certain threshold. Once this threshold is exceeded, the testing error, defined as the probability of whether the learner can understand a randomly selected text, drops sharply. Our analysis also provides the scaling law of the testing error with respect to the ratio $R$.
(ii) A semantic compression method: Two texts are said to be semantically equivalent if they require the same set of skills. A learner is considered generative if it can compose a semantically equivalent text for a set of learned skills. Once training for a generative learner is complete, the learner can be used as a semantic encoder/decoder. The idea is to index the learned skills and compress a text by using the indices of the skills required to understand that text. Since the testing error is very small when the threshold is exceeded, there is a very high probability that a randomly selected text can be compressed in this way. For texts not understood by the learner, we use traditional text compression methods.
@@
The paper is organized as follows: In Section II, we define a semantic language using a skill-text bipartite graph. In Section III, we define abstract learners, including 1-skill learners in Section III-A and Poisson learners in Section III-B. Density evolution analysis is employed to derive the scaling law for these abstract learners. In Section IV, we demonstrate how the trained learners can be utilized for semantic compression and semantic communication. The paper concludes in Section V, where we discuss possible extensions of our work.
@@ II-A Definition of a semantic language
Figure 1: An illustration for a skill-text bipartite graph, where the skill nodes connected to a text nodes indicates the set of skills required to understand the text. Two texts requiring the same set of skills are semnatically equivalent.
@@
We define that a skill $s$ in a semantic language is learned if the learner can determine whether the skill $s$ is present in any given text $t$. We consider a text $t$ to be understood by a learner if all the skills contained within $t$ are learned by the learner.
@@ II-B Learning a semantic language by sampling
(Uniform distribution) The probability that a skill $s$ is included in a randomly selected text $t$ is uniformly distributed, i.e., ${\bf\sf P}(s\in\phi(t))=1/|S|$.
From (A1)-(A2), the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|S|$ skill nodes on one side and $D$ text nodes on the other side. An edge exists between a skill node $s$ and a text node $t$ if $s\in\phi(t)$. The degree of a text node follows a Poisson distribution with mean $c$. According to (A1) and (A2), the number of edges between a text node and a particular skill node follows a Poisson distribution with mean $c/|S|$. Since there are $D$ text nodes, the number of edges of a skill node is also Poisson distributed, with mean $cD/|S|$. This leads to the following proposition.
@@ III Abstract learners
In this section, we develop a mathematical theory for learning a semantic language using abstract learners.
To a learner, all skills are novel before training. After presenting a text $t$ to the learner, a skill $s$ in the text $t$ might be learned. After repeatedly presenting the training set of $D$ texts, we aim to determine the fraction of skills that are learned by the learner.
III-A 1-skill learners
@@
(1-skill learner) An abstract learner is called a 1-skill learner if it can learn a novel skill $s$ by being presented with a text $t$ where the skill $s$ is the only novel skill in the text $t$. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other texts.
Now we describe how the 1-skill learner is trained. For each iteration, we present the $D$ sampled texts to the 1-skill learner in parallel. In the first iteration, a text containing only one skill is used to learn the skill in that text. Texts containing more than one skill do not contribute to learning in the first iteration. Once a skill $s$ is learned by the 1-skill learner, the number of novel skills in other texts containing $s$ is reduced by 1. In other words, the edges connected to the skill node $s$ can be removed from the skill-text bipartite graph. Therefore, the number of novel skills in the next iteration will not be greater than in the current iteration. In the second iteration, texts with only one novel skill are used to learn the skills in those texts. As in the first iteration, skills learned in the second iteration can be used to remove the corresponding edges in the skill-text bipartite graph. This iteration process is repeated until no more novel skills can be learned. We refer to this training process as Successive Cancellation of Novel Skills (SCNS).
This SCNS training process is mathematically equivalent to the iterative decoding approach in LDPC codes over the binary erasure channel (BEC) [15, 16, 17] and Irregular Repetition Slotted ALOHA (IRSA) [18, 19, 20, 21, 22, 23]. Following the density evolution analysis (in the asymptotic regime of a large number of skills) [24, 25, 26], let $q^{(i)}$ (respectively, $p^{(i)}$) be the probability that the skill end (respectively, the text end) of a randomly selected edge is not learned after the $i^{th}$ iteration of training. Initially, we have $q^{(0)}=1$. After the $i^{th}$ iteration, the number of novel skills in a randomly selected text follows a Poisson distribution with mean $q^{(i)}c$. For the 1-skill learner, a novel skill can be learned if it is the only novel skill in a text, which happens with probability $e^{-q^{(i)}c}$. Thus, we have
@@
Also, from SCNS, a skill end $s$ of a randomly selected edge is not learned after the $(i+1)^{th}$ iteration if the text nodes connected to $s$ (excluding the text end of the randomly selected edge) did not learn the skill $s$ after the $i^{th}$ iteration. Since the excess degree of a Poisson degree distribution is still Poisson with the same mean [11], the number of text nodes connected to $s$ (excluding the text end of the randomly selected edge) is a Poisson random variable with mean $cR$ (from Proposition 3(i)). Thus,
$q^{(i+1)}=\sum_{k=0}^{\infty}e^{-cR}\frac{(cR)^{k}}{k!}(1-p^{(i)})^{k}=e^{-cR(1-p^{(i)})}.$ (2)
@@
In view of (3), the testing error for the 1-skill learner is $1-p$, where $0\leq p\leq 1$ is the largest solution of (3). Since $p\geq 0$, we know from (4) that the testing error is at least $e^{-ce^{-cR}}$.
In Figure 2, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$. We compute $p$ by using the recursive equation $p^{(i+1)}=1-e^{-(ce^{-cR(1-p^{(i)})})}$ with the initial condition $p^{(0)}=1$. The number of iterations is set to 10,000. Since $1-p$ also represents the probability that a (test) text is understood by the learner, there exists a percolation threshold, indicated by a sudden change in the values of $p$. For $c=3$ (respectively, 4 and 5), this threshold occurs when $R$ exceeds approximately 1 (respectively, 1.5 and 2.5). The underlying rationale is that for $c\geq 3$, a randomly chosen text is unlikely to encompass precisely one novel skill. Consequently, a substantial volume of training texts is essential to ensure a sufficient quantity of texts initially containing only a single novel skill. This foundation of learned skills subsequently initiates a cascade of successive cancellations of novel skills.
@@
III-B Poisson learners
In this section, we extend the concept of the 1-skill learner to Poisson learners, where a novel skill can be learned with a certain probability.
@@ III-B Poisson learners
(Poisson learner) An abstract learner is termed a Poisson learner if a novel skill $s$ in a text $t$ can be learned with probability $P_{\rm suc}(\rho)$ when the number of novel skills in that text is Poisson distributed with mean $\rho$. Once a skill $s$ is learned, the learner is capable of identifying the skill $s$ in any other texts.
@@
Analogous to the argument for the 1-skill learner, the testing error for a Poisson learner is characterized in (4), where $0\leq p\leq 1$ is the largest solution in (7).
In Figure 3, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$ for the 2-skill learner. We compute $p$ by using the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-p)})$ with the initial condition $p^{(0)}=1$, where $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$. The number of iterations is set at 10,000. A percolation threshold, indicated by a sudden change in the values of $p$, exists for $c=4$ (respectively, 5) when $R$ is approximately 0.5 (resp. 0.6). The percolation thresholds for the 2-skill learner are significantly lower than those for the 1-skill learner, demonstrating that the 2-skill learner is a much faster learner than the 1-skill learner.

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated the emergence of capabilities (learned skills) when the number of system parameters and the size of training data surpass certain thresholds. The exact mechanisms behind such phenomena are not fully understood and remain a topic of active research. Inspired by the skill-text bipartite graph model proposed by Arora and Goyal for modeling semantic languages, we develop a mathematical theory to explain the emergence of learned skills, taking the learning (or training) process into account. Our approach models the learning process for skills in the skill-text bipartite graph as an iterative decoding process in Low-Density Parity Check (LDPC) codes and Irregular Repetition Slotted ALOHA (IRSA). Using density evolution analysis, we demonstrate the emergence of learned skills when the ratio of the size of training texts to the number of skills exceeds a certain threshold. Our analysis also yields a scaling law for testing errors relative to the size of training texts. Upon completion of the training, we propose a method for semantic compression and discuss its application in semantic communication.
@@ I Introduction
In the recent era of natural language processing (NLP), the evolution of large language models (LLMs), such as GPT-4 [1] and Gemini [2], has greatly impacted people’s daily lives. There is a growing consensus that enhancing language model performance and sample efficiency across a broad spectrum of downstream NLP tasks is closely linked to scaling up these models. This scaling involves increasing both the size of the training data and the number of model parameters, as discussed in [3, 4]. The relationship between the scale of an LLM and its performance can often be quantitatively forecasted using scaling laws, as explored in-depth in [5, 6].
@@
Recent studies [7, 8, 1] further demonstrates, through various numerical examples, that LLMs manifest emergent capabilities absent in their smaller counterparts. There is a notable enhancement in system performance once a certain critical scale threshold is surpassed, exhibiting a phase transition behavior often observed in network science [9]. However, as pointed out in [7], the exact mechanisms behind these emergent abilities in large language models are still not fully understood and remain a topic of active research.
@@
In the study presented by [10], a simple explanation is offered for the phase transition phenomenon observed in LLMs based on the concept of list decoders. The approach in [10] models an LLM as a sequence-to-sequence random function over a certain token space with $M$ possible tokens. It was shown that the expected number of erroneous sequences in the list decoder can be bounded by a constant if $M\epsilon<1$, where $\epsilon$ is the false alarm probability. Since transformer-based LLMs with more parameters and extensive training can memorize more patterns [11, 12], they are more likely to reduce the false alarm probability $\epsilon$ below the percolation threshold of $1/M$. However, the list decoder approach does not explain how the capabilities of language skills emerge from scaling up the size of the training data. In [13], a novel random graph approach was proposed to explain the scaling law discussed in [6]. A language is modeled by a random bipartite graph with skills on one side and texts (originally termed text-pieces in [13]) on the other side. An edge between a text and a skill indicates that understanding the text requires the skill. A text can be comprehended if all the skills connected to it are acquired.
Inspired by the skill-text bipartite graph model proposed in [13], we develop a mathematical theory to explain the emergence of learned skills by an abstract learner during the learning (or training) process. We summarize our contributions as follows:
  • (i)
@@
    Learning as an Iterative Decoding Process: We model the learning/training process for the skills in the skill-text bipartite graph as an iterative decoding process, similar to the analysis in Low-Density Parity Check (LDPC) codes [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. When a text is presented to a learner, a skill in that text might be learned with a certain probability (specific definitions for learners are given in Section III). By repeatedly presenting a large number of training texts to a learner, a fraction of skills can be learned.
  • (ii)
    The Emergence of Learned Skills: Following the density evolution analysis for the iterative decoding process (in the asymptotic regime of a large number of skills) [23, 24, 25], we show the emergence of learned skills when the ratio $R$ of the number of training texts to the number of skills exceeds a certain threshold. Once this threshold is exceeded, the testing error, defined as the probability of whether the learner can understand a randomly selected text, drops sharply. Our analysis also provides the scaling law of the testing error with respect to the ratio $R$.
  • (iii)
    The Association of Skills: Upon completing training, a learner acquires not just isolated skills but also understands how these skills interrelate, which can be represented by a skill association graph where edges connect learned skills appearing in the same text. Understanding this graph is crucial for making inferences, such as predicting the next skill from a given set. This process is akin to prompting a language model to forecast and generate text based on a skill set. This leads to a critical inquiry: Are the learned skills sufficiently interconnected to aid in prediction? To address this question, we use site (or node) percolation analysis [9] to derive conditions for the existence of a giant component in the skill association graph.
  • (iv)
    Hierarchy of Skills: We consider a setting with two classes of skills: basic and domain-specific. Learning a domain-specific skill first requires acquiring a random number of basic skills, establishing a hierarchy of skills. Initially, we train a foundation model on basic skills, which is then fine-tuned with domain-specific texts. Our findings identify two critical thresholds for minimizing testing errors in domain-specific tasks: (i) a large number of basic skills are learned when the number of basic texts exceeds the threshold in the foundation model, and (ii) a large number of domain-specific skills are learned when the number of domain-specific texts exceeds the threshold in the fine-tuning model.
  • (v)
    Application to Semantic Compression: We propose a semantic compression scheme that encodes the learned skills instead of all words in a text. In particular, we define the notion of a generative learner that is trained to identify the set of skills in a semantic language. Once trained, it can be used to perform semantic compression by representing a text using indices of the learned skills required to understand that text.
  • (vi)
    Multiple Classes of Skills and Texts: We extend the density evolution analysis to the settings with multiple classes of skills and texts, driven by the diversity of subjects in texts like math, physics, chemistry, law, etc. Such an extension is rooted on the framework of Poisson receivers with multiple classes of users and receivers in [26]. Interestingly, thresholds are also observed for the setting with multiple classes of skills and texts. The thresholds for various skill classes align due to the coupling of the density evolution equations, as illustrated in [26], resulting in the simultaneous emergence of various classes of skills.
@@
The paper is organized as follows: In Section II, we define a semantic language using a skill-text bipartite graph. In Section III, we introduce abstract learners, including 1-skill learners in Section III-A and Poisson learners in Section III-B. Density evolution analysis is employed to derive the scaling law for these abstract learners. In Section IV, we address the problem of learning the association between two learned skills and derive conditions for the existence of a giant component in the skill association graph. In Section V, we further address the problem setting with a hierarchy of skills by using a foundation model and a fine-tuning model. In Section VI, we demonstrate how the trained learners can be utilized for semantic compression and semantic communication. We present the extension to the setting with multiple classes of skills and texts in Section VII. The paper concludes in Section VIII, where we discuss possible extensions of our work.
@@ II-A Definition of a Semantic Language
Figure 1: An illustration for a skill-text bipartite graph, where the skill nodes connected to a text node indicates the set of skills required to understand the text. Two texts requiring the same set of skills are semantically equivalent.
@@
A skill $s$ in a semantic language is learned if the learner can determine whether the skill $s$ is present in any given text $t$ in $T$. A text $t$ is understood by a learner if all the skills contained within $t$ are learned by the learner.
@@ II-B Learning a Semantic Language by Sampling
(Uniform distribution) The probability that an edge of the (randomly selected) text node $t$ is connected to a skill node $s$ is $1/|S|$.
From (A1)-(A2), the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|S|$ skill nodes on one side and $D$ text nodes on the other side. An edge exists between a skill node $s$ and a text node $t$ if $s\in\phi(t)$. The degree of a text node follows a Poisson distribution with mean $c$. According to (A1) and (A2), the number of edges between a text node and a particular skill node follows a Poisson distribution with mean $c/|S|$. Since there are $D$ text nodes, the number of edges of a skill node is also Poisson distributed, with mean $cD/|S|$ (i.e., $cR$). This leads to the following proposition.
@@ III Abstract Learners
In this section, we develop a mathematical theory to understand the learning process of a semantic language by an abstract learner. To a learner, all skills are novel before training. After presenting a text $t$ to the learner, a skill $s$ in the text $t$ might be learned. After repeatedly presenting the training set of $D$ texts, we aim to determine the fraction of skills that are learned by the learner. We consider two types of learners based on their learning capabilities, namely, a 1-skill learner and a more general Poisson learner.
III-A 1-Skill Learner
@@
(1-Skill Learner) An abstract learner is called a 1-skill learner if it can learn a novel skill $s$ by being presented with a text $t$ where the skill $s$ is the only novel skill in the text. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other texts.
Now we describe how the 1-skill learner is trained. For each iteration, we present the $D$ sampled texts to the 1-skill learner in parallel. In the first iteration, a text containing only one skill is used to learn the skill in that text. Texts containing more than one skill do not contribute to learning in the first iteration. Once a skill $s$ is learned by the 1-skill learner, the number of novel skills in other texts containing $s$ is reduced by 1. In other words, the edges connected to the skill node $s$ can be removed from the skill-text bipartite graph. Therefore, the number of novel skills associated with each text in the next iteration will not be greater than that in the current iteration. In the second iteration, texts with only one novel skill remaining are used to learn the skills in those texts. As in the first iteration, skills learned in the second iteration can be used to remove the corresponding edges in the skill-text bipartite graph. This iteration process is repeated until no more novel skills can be learned. We refer to this training process as Successive Cancellation of Novel Skills (SCNS).
Figure 2: The Successive Cancellation of Novel Skills (SCNS) example.
In Figure 2, we present a simple example of the SCNS. Figure 2(a) depicts the original skill-text bipartite graph. Figure 2(b) focuses on the skill node $a$, which is connected to a text node with degree $1$; hence, we use a red edge to indicate that the skill node $a$ is learned by the text node. Consequently, the skill node $a$ can be learned, allowing the edges connected to it to be removed from the skill-text bipartite graph. Figure 2(c) shows the graph after the edges are removed, revealing that the skill node $b$ is connected to a text node with degree $1$, indicating it can also be learned. Figure 2(d) represents the last iteration, where no additional skills can be learned, concluding the process with the skills $a$ and $b$ being learned, while skills $c$ and $d$ are not.
This SCNS training process is mathematically equivalent to the iterative decoding approach in LDPC codes over the binary erasure channel (BEC) [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. Following the density evolution analysis (in the asymptotic regime of a large number of skills) [23, 24, 25], let $q^{(i)}$ be the probability that the skill end of a randomly selected edge is not learned after the $i^{th}$ iteration of training. Initially, we have $q^{(0)}=1$ (as all the skills are novel before training). After the $i^{th}$ iteration, the number of novel skills in a randomly selected text follows a Poisson distribution with mean $q^{(i)}c$ (a more detailed argument for this is provided in Section VII-A). For the 1-skill learner, a novel skill can be learned if it is the only novel skill in a text, which happens with probability $e^{-q^{(i)}c}$. Thus, the probability that the text end of a randomly selected edge is not learned after the $i^{th}$ iteration is
@@
Therefore, by SCNS, a skill end $s$ of a randomly selected edge can not be learned after the $(i+1)^{th}$ iteration if the text nodes connected to $s$ (excluding the text end of the randomly selected edge) were not able to learn the skill $s$ after the $i^{th}$ iteration. Since the excess degree of a Poisson degree distribution is still Poisson with the same mean [9], the number of text nodes connected to $s$ (excluding the text end of the randomly selected edge) is a Poisson random variable with mean $cR$ (from Proposition 3(i)). Thus,
$q^{(i+1)}=\sum_{k=0}^{\infty}e^{-cR}\frac{(cR)^{k}}{k!}{p^{(i)}}^{k}=e^{-cR(1-p^{(i)})}.$ (2)
@@
By the equality in (3), the testing error for the 1-skill learner (i.e., $1$ minus the probability in (4)) can be expressed as $p$, where $0\leq p\leq 1$ is the largest solution of (3). Since $p\geq 0$, we know from (4) that the testing error is at least $1-e^{-ce^{-cR}}$.
@@
In Figure 3, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$. We compute $p$ by using the recursive equation $p^{(i+1)}=1-e^{-(ce^{-cR(1-p^{(i)})})}$ with the initial condition $p^{(0)}=1$. The number of iterations is set to 10,000. Since $1-p$ also represents the probability that a (test) text is understood by the learner, there exists a percolation threshold, indicated by a sudden change in the values of $p$. The phase transition behavior is most evident when the average number of skills contained in each text (i.e., $c$) is greater than $3$. For $c=3$ (respectively, 4 and 5), this threshold occurs when $R$ exceeds approximately 1 (respectively, 1.5 and 2.5). The underlying rationale is that for $c\geq 3$, a randomly chosen text is unlikely to encompass precisely one novel skill. Consequently, a substantial volume of training texts is essential to ensure a sufficient quantity of texts initially containing only a single novel skill. This foundation of learned skills subsequently initiates a cascade of successive cancellations of novel skills.
III-B Poisson Learner
In this section, we extend the concept of the 1-skill learner to the Poisson learner, where multiple novel skills may be learned from a text but only with a certain probability.
@@ III-B Poisson Learner
(Poisson Learner) An abstract learner is termed a Poisson learner if a novel skill $s$ in a text $t$ can be learned with probability $P_{\rm suc}(\rho)$ when the number of novel skills in that text is Poisson distributed with mean $\rho$. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other text.
@@
Analogous to the argument for the 1-skill learner, the testing error for a Poisson learner is characterized in (4) with $p$ given in (7).